OpenStudy (anonymous):

which are the roots of 4x^2 + 8x + 3 = 0?

OpenStudy (anonymous):

\[-\frac{1}{2},-\frac{3}{2}\]

OpenStudy (anonymous):

sattelite73 can you help me out after

OpenStudy (anonymous):

maybe it isn't a conic section is it?

OpenStudy (anonymous):

btw this factors as \[(2x+1)(2x+3)=0\] which is why those are the solutions

OpenStudy (anonymous):

and where are you?

OpenStudy (anonymous):

the answer that my teacher told me is -1/2,-3/2 so?